Can you explain how to do this please? I get Z score of -4.44 but i dont think thats right because its not on the table
Find P(x < 63).
mean μ = 67 inches and standard deviation of σ = 3.6 inches. A random sample of 16 heights is obtained.

Answer by ewatrrr(24785)

 
Hi,
mean μ = 67 inches and standard deviation of σ = 3.6 inches. A random sample of 16 heights is obtained.
s = 3.6/sqrt(16) = .9
P(x<63) = P(z < (-4/.9))= P(z<-4.44) = 0.000004
Yes, that is quite a ways to the left :)
Below: z = 0, z = ± 1, z= ±2 , z= ±3 are plotted.
Note: z = 0 (x value the mean) 50% of the area under the curve is to the left and %50 to the right